Suppose the graph of a book stacking model has a slope of 4 and y-intercept of 0.
What is the thickness of each book?
Where did the stack begin?
What would be the height of the stack after 10 books?
Write a formula for the height of the stack after x books?
I am lost please help

Answer by stanbon(75887)

Suppose the graph of a book stacking model has a slope of 4 and y-intercept of 0.
T(b) = 4b + 0
-----------------------
What is the thickness of each book?
The slope describes the thickness of each book
T(1) = 4*1 = 4
Ans:: 4 units
-----------------------------
Where did the stack begin?
T(0) = 4*0 + 0 = 0
Ans:: 0
----------------------
What would be the height of the stack after 10 books?
T(10) = 4*10 + 0 = 40
Ans: 40 units
--------------------------
Write a formula for the height of the stack after x books?
T(x) = 4x
--------------
Cheers,
Stan H.
-----------------